OpenStudy (anonymous):

An Infinite Geometric sequence Such : \[S \infty =81/2 , Sn/Tn=(3^n-1)/2\] ,Find The G.s ???

OpenStudy (experimentx):

Sinf??

OpenStudy (anonymous):

Sum to infinity :D Its rule is S∞ =a/1-r .

OpenStudy (experimentx):

you should have written more explicitly

OpenStudy (anonymous):

what do u mean ?

OpenStudy (experimentx):

sum up to infinity

OpenStudy (anonymous):

An infinite geometric series is an infinite series whose successive terms have a common ratio. Such a series converges if and only if the absolute value of the common ratio is less than one ( | r | < 1 ).

OpenStudy (anonymous):

lol ,thats all what I know xD

OpenStudy (anonymous):

I think its we can apply the rule only ..

OpenStudy (experimentx):

put n=2 ...

OpenStudy (experimentx):

1/r + 1 = 8

OpenStudy (experimentx):

find the value of a from the left relation

OpenStudy (anonymous):

u mean 1-r ,but why u put a =1 ?

OpenStudy (experimentx):

Oh ... sorry ... forgot to divide by 2

OpenStudy (experimentx):

ignore that post ... i wasn't thinking

OpenStudy (anonymous):

nvm :)

OpenStudy (anonymous):

sry for asking many questions :/

OpenStudy (experimentx):

\[ \frac{S_2}{T_2} = 4 = \frac{a +ar}{ar}\] find the value of r from here get the value of a from other relation

OpenStudy (anonymous):

aha...

OpenStudy (anonymous):

@satellite73 : can u help :)

OpenStudy (anonymous):

@experimentX :from where u got that 4 in Sn/Tn=4=... ?

OpenStudy (experimentx):

from your second relation ... put n=2

OpenStudy (anonymous):

oh, OK

OpenStudy (experimentx):

you will get the value of r from here use your first relation to get the second value

OpenStudy (experimentx):

*get the value of a

OpenStudy (anonymous):

oh ,k........TY ^.^_^.^